OpenStudy (anonymous):

(6^7)^1/7 Simplify the expressions below. If your answer is not a whole number, enter it as a fraction in lowest terms, using the slash mark(/) as the fraction bar.

OpenStudy (anonymous):

\[(6^{7})^{1/7}\]

OpenStudy (jdoe0001):

\(\bf \LARGE a^{\frac{n}{m}} = \sqrt[m]{a^n}\)

OpenStudy (anonymous):

i dont quite understand.

OpenStudy (jdoe0001):

\(\bf (\Large {6^7)^{\frac{1}{7}} \implies \sqrt[7]{6^7}\\ \textrm{or one can say too}\\ (6^7)^{\frac{1}{7}} \implies (6^{\frac{7}{1}})^{\frac{1}{7}} \implies 6^{\frac{7}{1}\times\frac{1}{7}}}\)

OpenStudy (anonymous):

which is 6.

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (jdoe0001):

yw